[(45.5g)(0.0385 j/g 0k)( -373.0k )]+[(152g)(4.184j/g.k)(tf-291.7)]=0
1
Expert's answer
2018-01-29T06:24:36-0500
Solution:
[(45.5 g) × (0.0385 J/g⋅K) × (-373.0 K )] + [(152 g) × (4.184 J/g⋅K) × (Tf - 291.7 K)] = 0
-653.4 J + [(152 g) × (4.184 J/g⋅K) × (Tf - 291.7 K)] = 0
[(152 g) × (4.184 J/g⋅K) × (Tf - 291.7 K)] = 653.4 J
636 J/K × (Tf - 291.7 K)] = 653.4 J
Tf - 291.7 K = 1.03 K
Tf = 292.73 K